Sinking fund schedule

A ₱50,000 debt is to be repaid at the end of one year. The debtor establishes a sinking fund that earns 8% interest compounded quarterly. Construct a sinking fund schedule


1
Expert's answer
2021-05-19T17:50:31-0400

yearly effective rate(i)"=((1+\dfrac{r}{n})^n -1)\times 100"


r= periodic compounding rate

n= no. of periods or compounding frequency


yearly effective rate "(i)=((1+\dfrac{.08}{4})^4 -1)\times 100=8.243%" %


now we can calculate yearly payments as follows


"E(\dfrac{(1+i)^t – 1)}{i}) = FV"


Where


E=50000

i = yearly effective rate =8.243%

t= total period of fund=1 year

FV= Future Value of Fund


hence we have


"50000\times (\dfrac{(1.08243)^1 -1)}{0.08243}=FV"


solving this and we get 


FV=50000


now, the Sinking fund Schedule is as follows-


Yearly debt=50000

Interest at the end of year=8.234%

Future value of fund=50000